OpenStudy (anonymous):

Solve the exponential inequality. * Note: B5=base 5 logB5 x-7 is greater than or equal to -6

OpenStudy (anonymous):

@mathmale

OpenStudy (jdoe0001):

\(\Large \bf { log_5(x-7) \ge -6 \\ \quad \\ \textit{using the log cancellation rule of }{\color{blue}{ a^{log_ax}=x}}\qquad thus \\ \quad \\ log_5(x-7) \ge -6\implies {\color{blue}{ 5}}^{log_{\color{blue}{ 5}}(x-7)} \ge {\color{blue}{ 5}}^{-6}\\ \quad \\\implies x-7 \ge 5^{-6} }\) and I think you can take it from there